OpenStudy (anonymous):

How do I solve log(11x+9)=3+log(x+3)

OpenStudy (anonymous):

log a - log b = log(a/b)

myininaya (myininaya):

\[\log(11x+9)-\log(x+3)=3\] \[\log(\frac{11x+9}{x+3})=3\] \[\frac{11x+9}{x+3}=10^3\] \[11x+9=10^3*(x+3)\] \[11x+9=1000(x+3)\] here you should be able to solve from here